OpenStudy (anonymous):

P(x) with rational coefficients has the given roots (i and 7 + 8i) Find two additional roots?

OpenStudy (sirm3d):

Hint: the complex roots of a polynomial \(P(x)\) with rational coefficients come in pair as \(a+bi\) and \(a-bi\)

OpenStudy (anonymous):

I don't really understand that

OpenStudy (anonymous):

I know how to find roots but how do i go backwards to finding the equation given the roots?

OpenStudy (anonymous):

Oh do you mean the two roots would be put into form as: (x + i)(x - i) and (x + (7 +8i)) (x - (7 + 8i))?

OpenStudy (sirm3d):

close. one of the roots is \(i\) therefore its conjugate pair is \(-i\) the other root is \(7+8i\) so its conjugate pair is \(7-8i\)

OpenStudy (anonymous):

So (x + i)(x - i) isn't correct? I see how (x + (7 + 8i)) (x - (7 - 8i)) is correct but what do i do with the simplified root that get when i multiply?

OpenStudy (sirm3d):

the roots are \(i,\; -i,\; 7+8i,\;7-8i\). If you are reconstructing \(P(x)\), then \[P(x)=(x-i)(x-(-i))(x-(7+8i))(x-(7-8i))\]

OpenStudy (anonymous):

Ah ok. And once P(x) is known to find two additional roots use rational root theorem and synthetic division?

OpenStudy (sirm3d):

there is no need to use rational root and/or sythetic division for this problem. \(P(x)\) was reconstructed from the roots, so you knew the roots beforehand.

OpenStudy (anonymous):

The question i needed to answer originally though asked me for 2 additional roots given the roots i and 7 + 8i.

OpenStudy (anonymous):

The exact question is "A polynomial function P(x) with rational coefficients has the given roots. Find two additional roots of P(x)=0."

OpenStudy (sirm3d):

exactly. There is no need to write \(P(x)\). You only need to give the conjugate pair of each complex root. The conjugate pair of \(i\) is \(-i\), so you have \(-i\) as one of the two roots asked.

OpenStudy (anonymous):

Oh i see now... One last though, in the P(x) = .... you typed above in each parentheses it was always (x - the root or the roots opposite) is it always x MINUS? Is that just how its done?

OpenStudy (anonymous):

Oh and I'm sorry to keep asking questions separate from my original; but lets say one of the given roots is the square root of 10, would the inverse be i square root 10?

OpenStudy (sirm3d):

that's how reconstruction of \(P(x)\) is done. if one of the roots is \(\sqrt{10}\) which you can write as \(\sqrt{10}+0i\), its conjugate is \(\sqrt{10}-0i\)

OpenStudy (anonymous):

So after you get the reconstruction of P(x) and simplify it that's the final answer?

OpenStudy (sirm3d):

yes. but you only do that when \(P(x\) is asked.

OpenStudy (anonymous):

Well thank you very much for your help, i think i can figure these out now.

OpenStudy (sirm3d):

if you're up to it, you can write the polynomial \(P(x)\) with rational coefficients whose two roots are \(i\) and \(7+8i\).